Bolster Foods' (BF) balance sheet shows a total of $25 million long-term debt with a coupon rate of 8.50%. The yield to maturity on this debt is 8.00%, and the debt has a total current market value of $27 million. The balance sheet also shows that the company has 10 million shares of stock, and the stock has a book value per share of $5.00. The current stock price is $20.00 per share, and stockholders' required rate of return, rs, is 12.25%. The company
recently decided that its target capital structure should have 35% debt, with the balance being common equity. The tax rate is 40%. Calculate WACCs based on book, market, and target capital structures, and then find the sum of these three WACCs.

Please solve it by typing or hand, don't use EXCEL.
Answer #1

Solution:

In order to calculate WACC, we need to find the cost of debt, cost of equity and weights of the debt and equity.

Part A) WACC using book value

The cost of equity is given as 12.25%

The cost of debt based on books is 8.50% (Coupon)

The after tax cost of debt = 8.50*(1-Tax) = 8.5% * (1-0.40) = 5.1%

The debt value as per books = $25 million

The equity value based on books = Number of shares * book value per share = 10,000,000 * $5 = $50 million

Equity weight = Equity / (Equity + debt) = 50/(50+25) = 0.667

Debt weight = Debt / (Equity + debt) = 25/(50+25) = 0.333

WACC = Cost of equity * Weight of equity + Cost of debt * Weight of debt

             = 12.25% * 0.667 + 5.1% * 0.333 = 8.17%+1.70% = 9.87%

Part B) WACC using Market value

The cost of equity is given as 12.25%

The cost of debt based on market is 8 % (yield)

The after tax cost of debt = 8%*(1-Tax) = 8% * (1-0.40) = 4.8%

The debt value as per market = $27 million

The equity value based on Market = Number of shares * Market value per share = 10,000,000 * $20

= $2000 million

Equity weight = Equity / (Equity + debt) = 200/(200+27) = 0.881

Debt weight = Debt / (Equity + debt) = 27/(200+27) = 0.119

WACC = Cost of equity * Weight of equity + Cost of debt * Weight of debt

             = 12.25% * 0.881 + 4.8% * 0.119 = 10.79%+0.57% = 11.36%

Part C) WACC using Target debt equity ratio

The cost of equity is given as 12.25%

The cost of debt based on market is 8 % (yield)

The after tax cost of debt = 8%*(1-Tax) = 8% * (1-0.40) = 4.8%

Targeted debt weight = 0.35

Targeted equity weight = 0.65

WACC = Cost of equity * Weight of equity + Cost of debt * Weight of debt

             = 12.25% * 0.65 + 4.8% * 0.35 = 7.96% + 1.68% = 9.64%

Sum of these three WACC = 9.87% + 11.36% + 9.64% = 30.87%

Average of the WACC = 30.87% / 3 = 10.29%